Find the molecular formula of a compound that contains 18.7 g of lithium, 16.3 g of carbon, and 65.0 g of oxygen. The molar mass of the compound is 73.8 g/mol.
1
Expert's answer
2013-02-28T11:16:47-0500
Find the molecular formula of a compound that contains 18.7 g of lithium,
16.3 g of carbon, and 65.0 g of oxygen. The molar mass of the compound is
73.8 g/mol.
Answer:

The chemical formula for this compound is LixCyOz
We can find indexes from the following equation:

x : y : z = n(Li, moles) : n (C, moles) : n (O, moles) = [m (Li)/ MW(Li) ] : [ m (C) / MW (C) ] : [ m (O)/ MW (O) ] = [18.7/ 6.9] : [16.3/12.0] : [65.0/16.0] = 2.69 : 1.36 : 4.06& = 2 : 1: 3

The chemical formula for compound is& Li2CO3

Let's check the molar weight of compound :

MW (Li2CO3) = 2 x 6.9 + 12.0 + 3 x 16.0 = 73.8

Answer : Li2CO3